OpenStudy (lgbasallote):

an electric circuit contains an 8-ohm resistor in series with an inductor of 0.5 Henry and a battery of E volts. at t= 0, the current is 0. find E if the maximum current flowing in the circuit is 8 amperes

OpenStudy (anonymous):

in this question the resistor and the inductor are in a series connection, so at a time t->0('t' tending to zero(a limit)), the current in the circuit is zero. and at t=infinity the inductor looses its battle with the initial potential difference E and offers no back EMF(or impedance). so there is no longer any role of the inductor ,then E/R will be our max. current. we have E/R=8 and,\[R=8 \Omega\]so E=8x8=64V

OpenStudy (lgbasallote):

@SATYA.A.TEJA can you do this using differential equations? step-by-step?

OpenStudy (anonymous):

E=i(t)R+Ldi(t)/dt Solving the homogeneous equation: di(t)/i(t)=-(R/L)dt Then Log[i(t)]=-(R/L)t + K1 or i(t)=A1+A2*exp[-(R/L)t] Now apply boundary condition at t=0 i(0)=0----->A1+A2=0 and i(t)=A1*(1-exp[-(R/L)t]) Replacing i(t) in the differential equation, it turns out to be A1=E/R Then i(t)=E/R*(1-exp[-(R/L)t]) The maximum current happens when t->infinite and is E/R, then E=R*i(max)=8x8=64V

OpenStudy (lgbasallote):

is this the same as that equation \[\frac{dI}{dt} + (\frac{L}{R})I = \frac{E}{R}\]

OpenStudy (anonymous):

No, it is not the same. In fact it is: dI/dt+(R/L)I=E/L

OpenStudy (lgbasallote):

do you know how to use the formula i stated?

OpenStudy (anonymous):

The formula you stated is incorrect: Voltage in the coil = V1=LdI/dt Voltage in the resistor=V2=RI Voltage in the Battery=E It has to comply with Kirchoff's Law: E=V1+V2===>E=LdI/dt+RI (my formula, solved in my first answer)

OpenStudy (lgbasallote):

but this is what my teacher gave me...

OpenStudy (lgbasallote):

okay..can you just explain your formula step by step?

OpenStudy (lgbasallote):

just rechecked...it is R/L sorry so how to do this step by step?

OpenStudy (anonymous):

Here it goes. Enjoy it!

OpenStudy (lgbasallote):

thanks! will take a look